OpenStudy (anonymous):

Determine the value of k , such that the first polinomial is divisible by the second k^2x^4-3kx^2-4 ; x-1

OpenStudy (anonymous):

f(x) = 2(x2 - 3x) + 4 : factor 2 out in the first two terms = 2(x2 - 3x + (-3/2)2 - (-3/2)2) + 4 : add and subtract (-3/2)2 = 2(x - 3/2))2 + 17/2 : complete square and group like terms this a aid 2 finish u r task! hope this helps!! :-)

myininaya (myininaya):

x-1=0 x=1 so we want x=1 to be a zero of the polynomial which means we want f(1)=0 where f(x)=polynomial of course

OpenStudy (callisto):

Once you know the second binomial is a factor of first polynomial, By remainder theorem, f(1) = k^2(1)^4-3k(1)^2-4 =0 Then, you'll get k^2 - 3k -4 =0 Do the facorization for left hand side, you'll get (k-4)(k+1) =0 Solve the equation and you'll get 2 values for k.